# On the Moduli Space of Discrete Morse Functions

### v1

@@ 5. Merging two adjacent inner nodes:
Let $X$ be a regular CW complex, let $\sigma,\tau\in X$ be two 1 cells, and let $H_{\sigma}^{\tau}$ be the corresponding hyperplane in $\mathcal{H}$. Let $f\in\mathcal{M}$ be a discrete Morse function inside a region $R_{1}$ of $\mathcal{H}$ such that $\sigma$ and $\tau$ are critical and merge connected components, i.e. correspond to inner nodes of $M(f)$. Assume there is a region $R_{2}$ adjacent to $R_{1}$ via $H_{\sigma}^{\tau}\in\mathcal{H}$, i.e. there is a path from $R_{1}$ to $R_{2}$ that only crosses $H_{\sigma}^{\tau}$ and no other hyperplane of $\mathcal{H}$. If
@@
    $\sigma$ and $\tau$ correspond to adjacent inner nodes of $M(f)$, then moving through $H_{\sigma}^{\tau}$ first merges the corresponding nodes and then splits them the other way around.
@@
    the nodes corresponding to $\sigma$ and $\tau$ are not comparable in $M(f)$, passing through $H_{\sigma}^{\tau}$ corresponds to moving their associated values past each other without changing the underlying combinatorial merge tree.
@@
    If $\sigma$ and $\tau$ correspond to adjacent inner nodes of the induced merge tree, then $\sigma$ and $\tau$ subsequently merge three previously disconnected connected components into one. Crossing $H_{\sigma}^{\tau}$ means that the components are merged simultaneously while inside $H_{\sigma}^{\tau}$ and the merging order is reversed after exiting $H_{\sigma}^{\tau}$ on the other side. This corresponds by definition to first merging the nodes and then splitting them the other way around.
@@
Definition 5.15.
In order to topologize the space of merge trees $Mer$, we introduce the euclidean edit distance: We first define the metric for merge trees, whose underlying combinatorial merge trees are related by compositions of elementary edit morphisms (eem) in the same direction, i.e. either a composition of adding leaves and splitting inner nodes or removing leaves and merging inner nodes:
Let $\theta\in\operatorname{Map}_{\leq}(P,\mathbb{R}),\theta^{\prime}\in\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$ be merge trees such that $P$ and $P^{\prime}$ are related by a composition of eems in the same direction $\eta\colon P\rightarrow P^{\prime}$. Then by Remark 5.5 we have an induced morphism $\eta\colon P\rightarrow P^{\prime}$ if $\eta$ is a composition of adding leaves and splitting inner nodes. We extend the inclusion $\eta\colon P\rightarrow P^{\prime}$ associated to any composition of adding leaves and splitting inner nodes to a map $\eta_{*}\colon\operatorname{Map}_{\leq}(P,\mathbb{R})\times\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})\rightarrow\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$ by
(1) $\eta_{*}(\theta,\theta^{\prime})_{x}\coloneqq\begin{cases}\theta_{\eta^{-1}(x)}\text{ if }x\in\eta(P),\\ \theta_{\eta^{-1}(y)}\text{ if }x\notin\eta(P),\ x\text{ a leaf, }y\in\eta(P)\text{ is }x\text{'s parent in }P^{\prime}\\ \theta^{\prime}_{x}\text{ if }x\notin\eta(P)\text{ was added as a new parent node for a new leaf}\\ \theta^{\prime}_{y}\text{ if }x\notin\eta(P),\ x\text{ a leaf, }y\notin\eta(P)\text{ is }x\text{'s parent in }P^{\prime}\\ \theta^{\prime}_{y}\text{ if }x\notin\eta(P),\ x\text{ an inner node }y\text{ the node }x\text{ was split from}\end{cases}.$
If $\eta$ is a composition of removing leaves and merging inner nodes, we have an induced inclusion $\eta\colon P^{\prime}\rightarrow P$ in the other direction. Then we define
(2) $\operatorname{d}(\theta,\theta^{\prime})\coloneqq\operatorname{d}^{euc}(\eta_{*}(\theta,\theta^{\prime}),\theta^{\prime}).$
We define the euclidean edit distance by
(3) $\operatorname{d}(\theta,\theta^{\prime})\coloneqq\inf\limits_{\sigma}\{\sum\limits_{i}\operatorname{d}(\theta_{i},\theta_{i+1})\},$
where $\sigma=(\theta=\theta_{1},\dots,\theta_{n}=\theta^{\prime})$ are sequences of merge trees related by compositions of elementary edit morphisms in the same direction.
Proposition 5.16.
The euclidean edit distance $\operatorname{d}$ from Definition 5.15 is a pseudo metric on $Mer$ and a metric on $Mer_{<}$ and $Mer_{wb}$.
Proof.
The equality $\operatorname{d}(\theta,\theta)=0$ holds because the identity is an edit move. Symmetry holds because every edit move has an inverse: adding leaves is inverse to removing them and splitting inner nodes in inverse to merging them. Sequences of edit morphisms can be inverted stepwise by inverting elementary edit morphisms.
The triangle inequality follows from the triangle equality of the euclidean metric for trivial edit morphisms and by composition of the sequences in Equation 3 otherwise.
Let $\theta,\theta^{\prime}$ be merge trees over combinatorial merge trees $P,P^{\prime}$. If $P\cong P^{\prime}$, then positivity follows from the positivity of the euclidean metric. If $P\ncong P^{\prime}$ then $\operatorname{d}(\theta,\theta^{\prime})$ is by definition a sum of terms $\operatorname{d}(\theta_{i},\theta_{i+1})$ where $\theta_{i},\theta_{i+1}$ are merge trees over combinatorial merge trees $P_{i},P_{i+1}$ such that $P_{i}$ and $P_{i+1}$ are related by composition of elementary edit morphisms in the same direction. All these terms are non-negative because the euclidean metric is positive. Since $P\ncong P^{\prime}$, at least one of these terms must contain a non-trivial edit morphism. Let $\theta_{i},\theta_{i+1}$ be related by an non-trivial edit morphism. If $\theta_{i},\theta_{i+1}\in Mer_{<}$ then it follows from Equation 1 and Equation 2 that $\operatorname{d}(\theta_{i},\theta_{i+1})>0$ because values on new leaves/inner nodes coincide with the values of their parent nodes/nodes they are split from for $\eta_{*}(\theta_{i})$ but do not coinside for $\theta_{i+1}$. ∎
Theorem 5.17.
Let $X$ be a regular CW complex. Then the map $M\colon\mathcal{M}(X)\rightarrow Mer$ from Definition 5.8 that maps a discrete Morse function to its induced merge tree is continuous.
Proof.
Let $f_{0}\in\mathcal{M}(X)$ be an arbitrary discrete Morse function and let $\epsilon>0$. We have to show that there is a $\delta>0$ such that for all $f\in\mathcal{M}(X)$ with $\operatorname{d}(f,f_{0})<\delta$ we have $\operatorname{d}(M(f),M(f_{0}))<\epsilon$.
For that, we observe that for a fixed discrete Morse function $f_{0}$, the map to the induced merge tree is basically an orthogonal projection to the subspace of $\mathbb{R}^{X}$ which is spanned by the critical cells of dimensions 0 and 1 that correspond to the nodes of the induced merge tree. If $f_{0}$ lies in the interior of a region of $\mathcal{H}$, then we can choose $\delta$ small enough such that all $f$ with $\operatorname{d}(f,f_{0})<\delta$ lie in the same region of $\mathcal{H}(X)$ as $f_{0}$. Then all functions in the $\delta$ ball $U_{\delta}(f_{0})$ around $f_{0}$ induce the same underlying combinatorial merge tree and $M_{\lvert U_{\delta}(f_{0})}$ is just the orthogonal projection, which is continuous.
If $f_{0}$ lies on any of the hyperplanes of $\mathcal{H}$, then Lemma 5.10 and Lemma 5.13 describe how perturbations of $f_{0}$ affect the induced merge tree. That is, all merge trees in an $\epsilon$ neighborhood around $M(f_{0})$ that are induced by any discrete Morse function on $X$ are realized via a small perturbation of $f_{0}$. This is because merge trees in $U_{\epsilon}(f_{0})$ are related to $f_{0}$ by edit moves that correspond to crossing the hyperplanes that $f_{0}$ lies on. In this context, it is no problem that a priori $M(f_{0})$ admits more edit moves than $\mathcal{H}$ has hyperplanes because by Lemma 5.10 and Lemma 5.13, only the edit moves that correspond to a crossing a hyperplanes in the space of discrete Morse functions actually lead to merge trees that have preimages under the induced merge tree map. ∎
@@
Let $P$ be a poset of merge tree type. We define the leaf arrangement $\mathcal{A}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{A}(P)\coloneqq\{H_{x}^{y}|x\text{ is a leaf and }y\text{ is the parent node of }x\}$.
@@
Moreover, we define the braid arrangement $\mathcal{H}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{H}(P)\coloneqq\{H_{x}^{y}|x,y\in P\}$.
@@
    By Definition 5.8, the leaves of $P$ correspond to critical 0 cells of $f$. Moreover, the parent relation $x\prec y$ of any leaf $x$ corresponds to the collection of critical 1 cells that merge the connected component that corresponds to $x$, say $X_{M(f)_{y}-\varepsilon}[\sigma]$, where $\sigma$ is the unique critical 0 cell, with the connected components that correspond to $y$’s other child nodes. Let $\tau$ be one of these merging 1 cells. Then one of $\tau$’s boundary 0 cells $\sigma_{0}$ belongs to $X_{M(f)_{y}-\varepsilon}[x]$ and there is a gradient path $\gamma$ in $X_{M(f)_{y}}[\tau]$ that begins with $\tau\supset\sigma_{0}$. Since maximal gradient paths lead to critical cells and cannot lead through cells of increasing dimension, $\gamma$ has to lead to $\sigma$. If $f$ is weakly Morse–Benedetti, then each inner node $y$ corresponds to a unique merging 1 cell $\tau$ and because, by assumption, only one of $\tau$’s boundary 0 cells $\sigma_{0}$ belongs to $X_{M(f)_{y}-\varepsilon}[x]$, the gradient path starting at $\tau\sup\sigma_{0}$ is unique because it is contained in the 1 skeleton of $X$.
@@
    The hyperplanes of $\mathcal{O}(M(f)$ correspond by definition to inclusions of connected components sublevel complexes of $f$. Similar to the proof of (1), cover relations $x\prec y$ of nodes in $M(f)$ correspond to gradient paths from merging 1 cells that correspond to $y$ to any critical 0 cell of the connected component that corresponds to $x$. If $x\leq y$ is not a cover relation, then the statement follows inductively because the interval $[x,y]$ is a chain due to $P$ being of merge tree type.
@@
Since discrete gradient paths are sequences of matched cells, gradient paths correspond to intersections of hyperplanes of $\mathcal{A}(X)$. Even though this observation makes the correspondence between families of gradient paths and hyperplanes in Proposition 5.19 (1) into a correspondence between certain intersections of hyperplanes in $\mathcal{A}(X)$ and hyperplanes in $\mathcal{A}(M(f)$, that does not in general lead to a correspondence between cancellations of gradient paths and removals of leaves.
@@
If $f\colon X\rightarrow\mathbb{R}$ is a weak Morse–Benedetti function, then cancellations of critical 0 cells and 1 cells are in 1-1 correspondence with removals of leaves of the induced merge tree. If $f$ is not weakly Morse–Benedetti, cancellations of critical 0 cells and 1 cells still induce removals of leaves of the induced merge tree but not necessarily the other way around.
Corollary 5.22.
Let $P$ be a poset of merge tree type such that $P\cong M(f)$ for a weak Morse–Benedetti function $f\colon X\rightarrow\mathbb{R}$, then $M^{-1}(\mathcal{A}(P))\subset\mathcal{A}(X)$. Moreover, $M^{-1}(\mathcal{A}(P))$ agrees with $\mathcal{A}(X)$ on the 0 skeleton of $X$.

### v2

@@ 5. Merging two adjacent inner nodes:
Let $X$ be a regular CW complex, let $\sigma,\tau\in X$ be two 1 cells, and let $H_{\sigma}^{\tau}$ be the corresponding hyperplane in $\mathcal{H}$. Let $f\in\mathcal{M}$ be a discrete Morse function inside a region $R_{1}$ of $\mathcal{H}$ such that $\sigma$ and $\tau$ are critical and merge connected components, i.e. correspond to inner nodes of $M(f)$. Assume there is a region $R_{2}$ adjacent to $R_{1}$ via $H_{\sigma}^{\tau}\in\mathcal{H}$, i.e. there is a path from $R_{1}$ to $R_{2}$ that only crosses $H_{\sigma}^{\tau}$ and no other hyperplane of $\mathcal{H}$. Then the following statements hold:
@@
    If $\sigma$ and $\tau$ correspond to adjacent inner nodes of $M(f)$, then moving through $H_{\sigma}^{\tau}$ first merges the corresponding nodes and then splits them the other way around.
@@
    If the nodes corresponding to $\sigma$ and $\tau$ are not comparable in $M(f)$, passing through $H_{\sigma}^{\tau}$ corresponds to moving their associated values past each other without changing the underlying combinatorial merge tree.
@@
    If $\sigma$ and $\tau$ correspond to adjacent inner nodes of the induced merge tree, then $\sigma$ and $\tau$ subsequently merge three previously disconnected connected components into one. Crossing $H_{\sigma}^{\tau}$ means that the components are merged simultaneously while inside $H_{\sigma}^{\tau}$ and the merge order is reversed after exiting $H_{\sigma}^{\tau}$ on the other side. This corresponds by definition to first merge the nodes and then splitting them the other way around.
@@
[…]
@@
Moreover, we consider the braid arrangement $\mathcal{H}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$, which is given by $\mathcal{H}(P)\coloneqq\{H_{x}^{y}|x,y\in P\}$.
@@
    By Definition 4.26, the leaves of $P$ correspond to critical 0 cells of $f$. Moreover, the parent relation $x\prec y$ of any leaf $x$ corresponds to the collection of critical 1 cells that merge the connected component that corresponds to $x$, say $X_{M(f)_{y}-\varepsilon}[\sigma]$, where $\sigma$ is the unique critical 0 cell, with the connected components that correspond to $y$’s other child nodes. Let $\tau$ be one of these merging 1 cells. Then one of $\tau$’s boundary 0 cells $\sigma_{0}$ belongs to $X_{M(f)_{y}-\varepsilon}[x]$ and there is a gradient path $\gamma$ in $X_{M(f)_{y}}[\tau]$ that begins with $\tau\supset\sigma_{0}$. Since maximal gradient paths lead to critical cells and cannot lead through cells of increasing dimension, $\gamma$ has to lead to $\sigma$. If $f$ is weakly Morse–Benedetti, then each inner node $y$ corresponds to a unique merging 1 cell $\tau$ and because, by assumption, only one of $\tau$’s boundary 0 cells $\sigma_{0}$ belongs to $X_{M(f)_{y}-\varepsilon}[x]$, the gradient path starting at $\tau\supset\sigma_{0}$ is unique because it is contained in the 1 skeleton of $X$.
@@
    The hyperplanes of $\mathcal{O}(M(f))$ correspond by definition to inclusions of connected components sublevel complexes of $f$. Similar to the proof of (1), cover relations $x\prec y$ of nodes in $M(f)$ correspond to gradient paths, which lead from the merging 1 cells that correspond to $y$ to any critical 0 cell of the connected component that corresponds to $x$. If $x\leq y$ is not a cover relation, then the statement follows inductively because the interval $[x,y]$ is a chain due to $P$ being of merge tree type.
@@
Since discrete gradient paths are sequences of matched cells, gradient paths correspond to intersections of hyperplanes of the Morse arrangement $\mathcal{A}(X)$. Even though this observation makes the correspondence between families of gradient paths and hyperplanes in Proposition 4.35 (1) into a correspondence between certain intersections of hyperplanes in $\mathcal{A}(X)$ and hyperplanes in $\mathcal{A}(M(f))$, that does not in general lead to a correspondence between cancellations of gradient paths and removals of leaves.
@@
If $f\colon X\rightarrow\mathbb{R}$ is a weak Morse–Benedetti function, then cancellations of critical 0 cells and 1 cells are in 1-1 correspondence with removals of leaves of the induced merge tree. If $f$ is not weakly Morse–Benedetti, cancellations of critical 0 cells and 1 cells still induce removals of leaves of the induced merge tree but the canceled pair is not necessarily unique.
Definition 4.38.
Let $f$ be a discrete Morse function and let $\eta_{f}$ be a composition of splittings of inner nodes and additions of leaves. Then we denote by $\operatorname{Can}(\eta_{f})$ the collection of pairs of simplices whose cancellation via Corollary 4.37 corresponds to the removal of leaves that $\eta$ conducts. Moreover, we denote by $\operatorname{Mov}(\eta)$ the collection of critical 1-cells whose function values are changed in order to induce the splitting described by $\eta$, as is explained in Lemma 4.32.
Definition 4.39.
In order to topologize the space of merge trees $Mer$, we introduce the edit path distance: Let $\theta\colon P\rightarrow\mathbb{R},\theta^{\prime}\colon P^{\prime}\rightarrow\mathbb{R}$ be two merge trees and let $\gamma=(\eta_{n}\circ\dots\circ\eta_{1}\colon s(\theta)\rightarrow s(\theta^{\prime}))$ be a sequence of compositions of elementary edit moves in the same direction from $\theta$ to $\theta^{\prime}$. We define $\gamma_{0}\coloneqq\theta$, $\gamma_{n+1}\coloneqq\theta^{\prime}$, and $P_{i}\coloneqq\operatorname{im}(\eta_{i})$. Moreover, we consider sequences of height functions $\bar{\gamma}\coloneqq\gamma_{i}\colon P_{i}\rightarrow\mathbb{R}$. We define the length of $\gamma$ by
$\lambda(\gamma)\coloneqq\inf\limits_{\bar{\gamma}}\sum\limits_{i=0}^{n}\lvert\lvert{\eta_{i}}_{*}\gamma_{i}-\gamma_{i+1}\rvert\rvert$,
where $\lvert\lvert\cdot\rvert\rvert$ denotes the Euclidean norm.
We then define the edit path distance between $\theta$ and $\theta^{\prime}$ by
(2) $\operatorname{d}(\theta,\theta^{\prime})\coloneqq\inf\limits_{\gamma\colon s(\theta)\rightarrow s(\theta^{\prime})}\lambda(\gamma).$
Proposition 4.40.
The edit path distance $\operatorname{d}$ from Definition 4.39 is a pseudo-metric on $Mer$, $Mer_{<}$, and $Mer_{wb}$.
Proof.
The equality $\operatorname{d}(\theta,\theta)=0$ holds because the identity is an edit move. Symmetry holds because every edit move has an inverse: adding leaves is inverse to removing them and splitting inner nodes in inverse to merging them. Sequences of edit morphisms can be inverted stepwise by inverting elementary edit morphisms. The triangle inequality follows by construction as a path metric. Moreover, any two merge trees are connected by a sequence of edit moves, for example by merging all inner nodes and removing all leaves, which results in the trivial rooted tree. ∎
Remark 4.41.
The edit path distance cannot be a metric on $Mer$ because of the possibility of constant functions. These can result in paths of length 0 between merge trees that only differ in adjacent nodes with the same function value.
Moreover, the edit path distance cannot be a metric on $Mer_{<}$ and $Mer_{wb}$ either due to symmetries of the underlying combinatorial merge trees. It is possible to fix the latter issue by consideration of ordered underlying combinatorial merge trees, i.e. total order on the sets of children for each inner node. Then it is possible to fix representatives by demanding a suitable notion of compatibility between the orders on the sets of children and function values on the nodes.
Theorem 4.42.
Let $X$ be a regular CW complex. Then the map $M\colon\mathcal{M}(X)\rightarrow Mer$, where $\mathcal{M}(X)$ is equipped with the path distance [^18] See Definition 3.44. and Mer is equipped with the edit path distance, from Definition 4.26 that maps a discrete Morse function to its induced merge tree is Lipschitz continuous.
Proof.
We have to show that there is a constant $c\in\mathbb{R}$ such that for all $f,g\in\mathcal{M}(X)$ we have $\operatorname{d}(M(f),M(g))\leq c\cdot\operatorname{d}(f,g)$.
If $f,g$ lie in the interior of the same region $R$ of $\mathcal{H}$, then $f$ and $g$ have the same critical cells, the same gradient field, and induce the same combinatorial merge tree, so $M$ maps $f$ and $g$ to the same subspace, so $M_{\lvert R}$ is just an orthogonal projection, which is Lipschitz continuous with $c=1$. In this case, the path metric and the edit path distance both reduce to the Euclidean metric.
For the case that $f,g$ lie in different regions $f\in R_{1},g\in R_{2}$ of $\mathcal{H}$, we consider that by Proposition 4.35, Lemma 4.29, Lemma 4.32, and Corollary 4.37 the necessary edit morphisms $\tilde{\gamma}$ between the induced combinatorial merge trees correspond to paths $\gamma$ in the space of discrete Morse functions that cross certain hyperplanes $H_{i}$ of $\mathcal{H}$.
Let $\tilde{\gamma}$ be any sequence of edit moves from $s(M(f))$ to $s(M(g))$ and let $\gamma$ be a corresponding path in the space of discrete Morse functions from $f$ to $g$. We will prove the statement of the theorem by showing that $)\lambda(\gamma)\leq(\lambda(\tilde{\gamma}$)).
Let $(\tilde{\gamma}_{i})$ be a sequence of intermediate height functions as in Definition 4.39. Due to each of the $\tilde{\gamma}_{i}$ lying inside an intersection of hyperplanes of the braid arrangement on $Mer$ (or in a face reached by exiting a certain intersection of hyperplanes), the corresponding path $\gamma$ has a corresponding subdivision with intermediate functions $h_{i}$ inside the corresponding intersections of hyperplanes of the braid arrangement (or in a face reached by exiting the corresponding intersection of hyperplanes) on the space of discrete Morse functions. The nodes of the induced merge tree only correspond to certain critical 1- and all critical 0-cells of $X$, which implies for the computation of the length of path segments in the space of merge trees, that the relevant vectors can be embedded as subvectors in the relevant vectors in the space of discrete Morse functions. Hence, we have $\lvert\lvert{\eta_{i}}_{*}\gamma_{i}-\gamma_{i+1}\rvert\rvert\leq\lvert\lvert h_{i}-h_{i+1}$ for all $i$ even in the space of discrete functions. Moreover, the path $\gamma$ may have to cross additional intersections of hyperplanes in order to stay within the space of discrete Morse functions, which could only potentially increase the length of $\gamma$ due to the triangle inequality.
Since this inequality between the lengths of path segments holds for all segments of each potential path, it also holds for the respective infima and, hence, also for the distance $d(M(f),M(g)\leq d(f,g)$. Hence, the induced merge tree $M$ satisfies Lipschitz-continuity with $c=1$.
Remark 4.43.
Lipschitz continuity of invariants of filtered spaces is an especially relevant property for applications in TDA. In this framework, being Lipschitz continuous is also called satisfying stability and is a necessary property for invariants to be used for statistical analysis. Stability of merge trees is by no means a new result, for example in [CCL+22], the authors have proved stability of the induced merge tree with respect to a large class of distances on merge trees.
Example 4.44.
To illustrate the ideas of the proof, we give an example of the distance between two discrete Morse functions on a regular CW complex and their induced merge trees:
[…]
On the side of the induced merge trees, we describe the necessary edit moves from $M(f)$ to $M(g)$: First, we merge the lime-colored inner node labeled 3 with the pink-colored inner node labeled 4. Then we split off a new inner node from the pink node the other way around, label it 3, and merge the orange node with new newly created node. Throughout the whole process, we assign the green node labeled 1 as a child to the inner node we move around. As a result, the fixed nodes have distance zero and the distance between the merge trees is purely given by the edit costs.
$d(M(f),M(g))=1+1+1=3.$
Regarding the distance between the discrete Morse functions, we consider that there are 6 cells whose function values with respect to $f$ and $g$ lie in different faces of $\mathcal{H}$: all 1 cells except for the one labeled 5, and the two 2 cells of the sphere on the left. We denote them as follows:
[…]
We consider the relevant differences in function values that create the necessity of crossing certain hyperplanes:
  1. (1)
    $(\alpha)>f(\beta),\ g(\alpha)=g(\beta)$
  2. (2)
    $f(\alpha)<f(\epsilon),\ g(\alpha)>g(\epsilon)$
  3. (3)
    $f(\beta)=f(\gamma),\ g(\beta)>g(\gamma)$
  4. (4)
    $f(\beta)<f(\epsilon),\ g(\beta)>g(\epsilon)$
  5. (5)
    $f(\beta)<f(\rho),\ g(\beta)=g(\rho)$
  6. (6)
    $f(\gamma)<f(\epsilon),\ g(\gamma)=g(\epsilon)$
  7. (7)
    $f(\delta)=f(\epsilon),\ g(\delta)>g(\epsilon)$
  8. (8)
    $f(\epsilon)>f(\rho),\ g(\epsilon)<g(\rho)$
It follows that in order to get from $f$ to $g$ in the space of discrete Morse functions, at least 11 entries/exits of hyperplanes need to happen, which result in 11 crossing costs. The crossings related to the changes in (8) and (6) correspond to the above described edit moves on the induced merge trees and have the same cost. The other necessary crossings add to the distance between the functions but are not reflected in the distance between the induced merge tree.
The optimal path from $f$ to $g$ turns out to be given by the direct line segment between $f$ and $g$. If we parametrize it by $\gamma(t)\coloneqq(1-t)\cdot f+t\cdot g$, the hyperplanes that correspond to (3) and (7) are exited at $t=0$, the hyperplanes that correspond to (2) and (8) are being crossed at $t=\frac{1}{2}$, the hyperplane that corresponds to (4) is being crossed at $t=\frac{2}{3}$, and the hyperplanes that correspond to (1), (5), and (6) are entered at $t=1$. The crossing costs are minimized by taking the cut points in the middle between the crossings. This results in a distance of
$d(f,g)=\frac{3}{4}+\frac{3}{4}+\frac{1}{4}+\frac{1}{4}+\frac{1}{2}+\frac{1}{2}=3$
Since in this example all intersections of hyperplanes that are being crossed by the direct line between $f$ and $g$ are actually contained in the space of discrete Morse functions and the function values agree on matched cells, the distance agrees with the one between the induced merge trees.